OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (tyteen4a03):

Let sin(x) be y, solve the quadratic equation in terms of y then find x.

OpenStudy (anonymous):

could you work it out please thank you

OpenStudy (tyteen4a03):

@lebronjames We're not here to do your homework. You'll have to figure it out by yourself. You can however ask questions when you're stuck on a particular step.

OpenStudy (anonymous):

im stuck on what to do thats why im asking

OpenStudy (tyteen4a03):

@lebronjames Let sin(x) be y, then the equation sin2x + sin x = 0 would become y^2 + y = 0.

OpenStudy (anonymous):

then y^2 = y and y = sqrtY ?

OpenStudy (anonymous):

@lebronjames what could would it to do have y=sqrt(y)?

OpenStudy (anonymous):

What good*

OpenStudy (anonymous):

nothing

OpenStudy (tyteen4a03):

@lebronjames Use the quadratic equation to get the solutions to this problem. (If you have no idea what they are, you may need to retake Algebra II)